ABCD is a trapezium. AB is parllel to DC and AO/CO=BO/DO. then find the value of x if 
OA=2x+7, OB=4x, OD=4x-4 and OC=2x+4

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
From Given conditionAO/CO=BO/DO
By substituting the values given in above expression we get
(2x+7)/(2x+4)=(4x)/(4x-4)
By cross multiplication we get
(2x+7)(4x-4)=(4x)(2x+4)
8x2-8x+28x-28=8x2+16x
On simplification we get x=7